OpenStudy (bananas1234):

Simplify completely: square root of 81y^5

OpenStudy (anonymous):

59049

OpenStudy (bananas1234):

i dont think that is what i need sorry.

OpenStudy (bananas1234):

options: A. 9y^2 square root y B. 3y^2 C. 3y^2 square root 2y D. 9y square root y

OpenStudy (bananas1234):

ok

OpenStudy (bananas1234):

why?

OpenStudy (bananas1234):

ok i think that makes sense

OpenStudy (bananas1234):

ok thanks!

pooja195 (pooja195):

Whats the square root of 81?

OpenStudy (bananas1234):

9

pooja195 (pooja195):

good \[\huge~9*(y*y*y*y*y)\]

OpenStudy (anonymous):

\[\sqrt{81y^{5}} = \sqrt{81}*\sqrt{y^{4}}*\sqrt{y}\] You can think of it like this, which seems to be the approach of pooja above. Good luck :)

OpenStudy (bananas1234):

9y

OpenStudy (bananas1234):

9y square root y?

OpenStudy (bananas1234):

@pooja195

pooja195 (pooja195):

I agree with that. @Concentrationalizing is it right though? :/

OpenStudy (anonymous):

Well, given \[\sqrt{81y^{5}} = \sqrt{81} * \sqrt{y^{4}} * \sqrt{y}\] \(\sqrt{81} = 9\) For the \(\sqrt{y^{4}}\) part, if we think of the square root as being an exponent of 1/2, we have: \[\sqrt{y^{4}} = y^{4*\frac{1}{2}} = y^{2}\] The \(\sqrt{y}\) part cannot be simplified. Put it all together and we have \(9y^{2}\sqrt{y}\)

OpenStudy (bananas1234):

i think i understand that, Thank you!

OpenStudy (anonymous):

Of course :)